# Echoes of the past: A unified perspective on fading memory and echo states

### v1

@@ 1.  Introduction
Recurrent neural networks (RNNs) have become increasingly popular as machine learning models in tasks involving time series and temporal data generated by dynamical systems. What makes RNNs so popular in applications is their efficient scalability and ease of training, owing to model designs such as reservoir computing (RC). Other machine learning paradigms such as large language models require vast amounts of training data and cost-intensive computational resources for the initial training as well as retraining sessions, which brings increasingly difficult challenges in its wake pertaining to energy consumption and cost of chatbot queries. [1, 13]. In contrast, RC is designed so that only a small part of the model is being trained algorithmically, the main part of the model being chosen randomly [12, 19] or even realized by a physical system. The latter has sparked a particular interest in the physics and robotics community, which explore different systems that can function as these reservoirs. The first example of such a physical reservoir was a bucket of water, whose computational ability resides in translating input sound waves to ripples of waves on the surface of the water [5]. Since then many other physical reservoirs have been proposed from photonics and analog circuits to mechanical and biological bodies, to name only a few [26, 30]. A particularly promising direction is the use of quantum systems as reservoirs for so-called quantum RC [6, 7, 25, 8].
@@
On the theoretical side, a fundamental property of RNNs is their ability to create reliable input/output responses. This ability is often linked to how the network handles its memory of the information it processed – or rather forgets its memory. Over the past decades, different notions have been brought forth to conceptualize the decay of its memory. Prominent keywords are steady states [4, 2], echo states, state forgetting, input forgetting [12], and fading memory [2]. It is a long-standing folklore belief that these notions can be used interchangeably. Although some subtleties concerning this belief have been discovered in recent years [27, 29], a deeper understanding of the relationship between these various notions is lacking. It is difficult to analyze the precise relationships in part because the notions are not even defined using the same terminology. This gap will be filled in the present work: we unify all notions and present new results linking them.
@@
Broadly speaking, an RNN processes a sequence of inputs $(\underline{u}_{t})_{t}$ and produces a sequence of states $(\underline{x}_{t})_{t}$, where at time $t$ the current input $\underline{u}_{t}$ and the previous state $\underline{x}_{t-1}$ are transformed into the next state $\underline{x}_{t}$ by the so-called state map $f$. In applications, it is a crucial part of the model design that the states are further transformed by a readout map to yield the final output. But adding a readout does not affect the underlying dynamics or its memory of the processed information and is not of relevance in this work. When an RNN ought to create a reliable response to a given input sequence, the state $\underline{x}_{t}$ at time $t$ is expected to depend to much greater extent on the most recent inputs $\underline{u}_{t},\underline{u}_{t-1},\dots$ than on inputs $\underline{u}_{s}$, s<<ts<\!\!<t, from the distant past. This naturally leads to the idea that the system is ‘input forgetting’ [12, 10]. The states, in this case, have been coined ‘steady states’ or ‘steady state solution’ in earlier works [4, 2] and ‘echo states’ in later works [12]. We caution that sometimes the terminology ‘echo states’ includes the requirement that the sequence of states is unique for each given input sequence.
@@
The same idea as for input forgetting lies behind the notion of ‘fading memory’ but the mathematical formulation differs. Customarily, fading memory is defined as continuous dependence of the echo states on the input sequence [2, 10]. On first inspection, it may not be clear how continuous dependence captures the idea of fading memory. The key to this definition is the choice of topology with respect to which continuity is required. Indeed, some topologies on the space of input sequences boast the property that two sequences are close if their recent entries are close even if their distant past entries differ significantly. Since there are several topologies with this property, there are different notions of fading memory depending on the choice of topology. Although they are closely related to each other and sometimes overlap, various choices give rise to non-identical notions of fading memory [27, 29]. The notions of input forgetting and fading memory are concerned with echo states of distinct input sequences that are similar to each other in the recent past. On the contrary, the notion of ‘state forgetting’ considers a single input sequence and two different initial states. If the next states produced by the state map become increasingly similar to each other as time evolves regardless of the dissimilarity of the initial states, then the system is state forgetting [12, 10]. The initial states may not be part of the sequence forming the input’s echo states, but some form of attraction is at play, where the next states obtained from the initial ones approach the echo states. Indeed, one of the overarching principles governing all these various notions is that the echo states form the dynamical attractor of the RNN to which the states converge [12, 3, 22, 29].
@@
In this work, we unify the various notions surveyed above in a common language, in which it will become easier to see how they are linked to forward and pullback attraction of the echo states. We will derive new implications and equivalences between several notions, and along the way we will encounter alternative and shorter proofs to some of the classical results in the literature. The main notions and statements are discussed in Section˜2, and the proofs together with additional details are presented in Section˜2.6.
@@ 2.3  State and input forgetting
Figure 1: Relations between state and input forgetting properties.
@@
$\mathcal{R}=\{x\in\mathcal{X}\colon\text{there exists }(\underline{x},\underline{u})\in\mathcal{S}\text{ with }\underline{x}_{0}=x\}.$
@@
From the given definition, it may not be immediately apparent why the IFP indeed conceptualizes ‘input forgetting’. To see how the IFP is motivated, consider $x^{1},x^{2}\in\mathcal{R}$. Take solutions $(\underline{x}^{1},\underline{u}^{1}),(\underline{x}^{2},\underline{u}^{2})\in\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}^{-}$ with $\underline{x}_{0}^{1}=x^{1}$ and $\underline{x}_{0}^{2}=x^{2}$. Then, $\psi_{n}(x^{i},\underline{u})$ is exactly the time zero component of a solution for the input $\underline{u}^{i,n}:=(\dots,\underline{u}_{-1}^{i},\underline{u}_{0}^{i},\underline{u}_{1},\dots,\underline{u}_{n})$. The $n$ most recent inputs in the sequences $\underline{u}^{1,n}$ and $\underline{u}^{2,n}$ are the same. Thus, although the IFP is defined through states $x^{1},x^{2}\in\mathcal{R}$, it really considers solutions of different input sequences whose difference lies further and further in the past as $n$ grows.
@@
The implications and equivalences depicted in the diagram in Figure˜1 hold by definition.
@@ 2.4  Attraction
We hinted in the introduction that one of the overarching principles is that the echo states form the dynamical attractor of the RNN to which the states converge. To make this claim rigorous, let us specify the precise dynamical system. Despite being intrinsically non-autonomous, it was proposed in [29] to model the dynamics autonomously on an extended sequence space. To this end, consider the map $\varphi\colon\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}\rightarrow\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}$ given by $\varphi(\underline{x},\underline{u})=((\underline{x},f(\underline{x}_{0},\underline{u}_{1})),\sigma(\underline{u}))$, whose iterates can be seen as extensions of the maps $\psi_{n}$. Concretely, $\psi_{n}\circ(p_{0}\times\pi)=p_{0}\circ\varphi^{n}$, where $p_{0}\colon\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}\rightarrow\mathcal{X}$ and $\pi\colon\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}\rightarrow\underline{\mathcal{U}}$ are the projections $p_{0}(\underline{x},\underline{u})=\underline{x}_{0}$ and $\pi(\underline{x},\underline{u})=\underline{u}$. This relation implies that $\psi_{n}(x,\underline{u})$ belongs to the $p_{0}$-projection of the fiber $\pi^{-1}(\sigma^{n}(\underline{u}))$, and $\psi_{n}(x,T^{n}(\underline{u}))$ belongs to the $p_{0}$-projection of the fiber $\pi^{-1}(\underline{u})$. It has been shown in [29] that the global attractor $\mathcal{A}:=\bigcap_{n\in\mathbb{N}_{0}}\varphi^{n}(\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}})$ of $\varphi$ is exactly equal to $\tau^{-1}(\mathcal{S})$. This implies that if $x\in\mathcal{R}$, then $\psi_{n}(x,\underline{u})$ and $\psi_{n}(x,T^{n}(\underline{u}))$ belong to the $p_{0}$-projection of the attractor $\mathcal{A}$. The fibers $\mathcal{A}\cap\pi^{-1}(\underline{u})$ of the attractor capture pullback attraction, and the fibers $\mathcal{A}\cap\pi^{-1}(\sigma^{n}(\underline{u}))$ capture forward attraction. We now see that the SFP realizes forward attraction: if the SFP holds, then for all $\underline{u}\in\underline{\mathcal{U}}$ and $x\in\mathcal{X}$
$\begin{split}&\limsup_{n\rightarrow\infty}\mathrm{dist}(\psi_{n}(x,\underline{u}),\mathcal{A}\cap\pi^{-1}(\sigma^{n}(\underline{u})))\\ &\leq\limsup_{n\rightarrow\infty}d_{\mathcal{X}}(\psi_{n}(x,\underline{u}),\psi_{n}(x^{\prime},\underline{u}))=0,\end{split}$
where $x^{\prime}$ is an arbitrary element in $\mathcal{R}$. Likewise, the s-SFP realizes pullback attraction:
@@
$\begin{split}&\limsup_{n\rightarrow\infty}\mathrm{dist}(\psi_{n}(x,T^{n}(\underline{u})),\mathcal{A}\cap\pi^{-1}(\underline{u}))\\ &\leq\limsup_{n\rightarrow\infty}d_{\mathcal{X}}(\psi_{n}(x,T^{n}(\underline{u})),\psi_{n}(x^{\prime},T^{n}(\underline{u})))=0.\end{split}$
@@
Figure 2: Relations between state and input forgetting properties in Theorem˜2.4. The blue implication arrows hold if $\underline{\mathcal{X}}^{-}$ is compact; the green implication arrows hold if $\underline{\mathcal{X}}^{-}$ and $\underline{\mathcal{U}}$ are compact; the purple implication arrows do not hold in general.
2.5  Main result
In our main result, stated below, we extend the diagram in Remark˜2.2 with additional implication arrows under compactness assumptions. It generalizes [12, Proposition 1], [21, Theorem 1], and [18, Remark 1], and provides alternative and shorter proofs.
Theorem 2.4.
Consider the diagram depicted in Figure˜2. If $\underline{\mathcal{X}}^{-}$ is compact, then the implication arrows indicated in blue are valid; and if $\underline{\mathcal{X}}^{-}$ and $\underline{\mathcal{U}}$ are compact, then the implication arrows indicated in green are valid. Furthermore, we will see counterexamples to the implication arrows indicated in purple.

### v2

@@ 1.  Introduction
Recurrent neural networks (RNNs) have become increasingly popular as machine learning models in tasks involving time series and temporal data generated by dynamical systems. What makes RNNs so popular in applications is their efficient scalability and ease of training, owing to model designs such as reservoir computing (RC). Other machine learning paradigms such as large language models require vast amounts of training data and cost-intensive computational resources for the initial training as well as retraining sessions, which brings increasingly difficult challenges in its wake pertaining to energy consumption and cost of chatbot queries. [5, 24]. In contrast, RC is designed so that only a small part of the model is being trained algorithmically, the main part of the model being chosen randomly [23, 35] or even realized by a physical system. The latter has sparked a particular interest in the physics and robotics community, which explore different systems that can function as these reservoirs. The first example of such a physical reservoir was a bucket of water, whose computational ability resides in translating input sound waves to ripples of waves on the surface of the water [12]. Since then many other physical reservoirs have been proposed from photonics and analog circuits to mechanical and biological bodies, to name only a few [44, 54]. A particularly promising direction is the use of quantum systems as reservoirs for so-called quantum RC [13, 14, 43, 39, 40, 15].
@@
On the theoretical side, a fundamental property of RNNs is their ability to create reliable input/output responses. This ability is often linked to how the network handles the information it processes – or rather, forgets it. Over the past few decades, various notions have been proposed to conceptualize the decay of its memory. Prominent keywords are steady states [8, 6], echo states, state forgetting, input forgetting [23], and fading memory [6]. It is a long-standing folklore belief that these notions can be used interchangeably. Although some subtleties concerning this belief have been identified in recent years [36, 45, 47], a deeper understanding of the relationship among these notions is lacking. Such a task is difficult in part because the notions are not even defined using the same terminology. This gap will be filled in the present work, in which we unify all notions and present new results linking them.
@@
Broadly speaking, an RNN processes a sequence of inputs $(\underline{u}_{t})_{t}$ and produces a sequence of states $(\underline{x}_{t})_{t}$, where at time $t$ the current input $\underline{u}_{t}$ and the previous state $\underline{x}_{t-1}$ are transformed into the next state $\underline{x}_{t}$ by the so-called state map $f$. In applications, it is a crucial part of the model design that the states are further transformed by a readout map to yield the final output. But adding a readout does not affect the underlying dynamics or its memory of the processed information and is not of relevance in this work.
There is a large spectrum of modeling situations in which, in the words of Volterra [56], an “extremely natural postulate is to suppose that the influence of the (input) a long time before the given moment gradually fades out". Such circumstances are pervasive in physical systems with friction or dissipation and in econometrics, where most parametric models are designed to exhibit some sort of fading memory-type feature [10, 20]. Similar ideas appear in cybernetics, where adaptation to a changing environment requires feedback controllers to rely on current signals rather than outdated ones [3, 57]. This same principle underlies modern adaptive algorithms in computer science: in online learning, estimators employ decaying step sizes or explicit forgetting mechanisms to remain responsive to non-stationary data streams [51]; reinforcement learning algorithms impose exponential decay through temporal-difference methods and eligibility traces [53]; and streaming algorithms and adaptive filters downweight older observations to efficiently track evolving patterns [9, 22]. Across these fields, emphasizing recent information is not merely a heuristic but a formally justified response to environments in which the data-generating process itself changes over time.
This emphasis on recent information is also well established and standard practice in control and system-identification [50], particularly when the underlying system is subject to time variation or unmodeled disturbances. Optimal state estimators such as the Kalman filter assign greater weight to new measurements through the Kalman gain, which increases when model uncertainty grows and thus emphasizes the informational value of recent data [41, 52]. Adaptive control and recursive parameter estimation use exponential forgetting factors and finite-memory estimators to enable effective tracking of systems whose characteristics drift over time [4, 33].
In all the situations that we just described, one is naturally led to systems that are ‘input forgetting’ [23, 19]. The states, in this case, have been coined ‘steady states’ or ‘steady state solution’ in earlier works [8, 6] and ‘echo states’ in later works [23]. We caution that sometimes the terminology ‘echo states’ includes the requirement that the sequence of states is unique for each given input sequence.
@@
Even though the same modeling intuition underlies the notions of ‘input forgetting’ and ‘fading memory’, their mathematical formulations differ. Customarily, fading memory is defined as continuous dependence of the echo states on the input sequence [6, 19]. On first inspection, it may not be clear how continuous dependence captures the idea of fading memory. The key to this definition is the choice of topology with respect to which continuity is required. Indeed, some topologies on the space of input sequences boast the property that two sequences are close if their recent entries are close even if their distant past entries differ significantly. Since there are several topologies with this property, there are different notions of fading memory depending on the choice of topology. Although they are closely related to each other and sometimes overlap, various choices give rise to non-identical notions of fading memory [45, 47]. The notions of input forgetting and fading memory are concerned with echo states of distinct input sequences that are similar to each other in the recent past. On the contrary, the notion of ‘state forgetting’ considers a single input sequence and two different initial states. If the next states produced by the state map become increasingly similar to each other as time evolves regardless of the dissimilarity of the initial states, then the system is state forgetting [23, 19]. The initial states may not be part of the sequence forming the input’s echo states, but some form of attraction is at play, where the next states obtained from the initial ones approach the echo states. Indeed, one of the overarching principles governing all these various notions is that the echo states form the dynamical attractor of the RNN to which the states converge [23, 7, 38, 47].
@@
In this work, we unify the various notions surveyed above in a common language, in which it will become easier to see how they are linked to forward and pullback attraction of the echo states. We will derive new implications and equivalences between several notions, and along the way we will encounter alternative and shorter proofs to some of the classical results in the literature. The main notions and statements are discussed in Section˜2, and they will be related to notions from the literature in Section˜3.
@@ 2.3  State and input forgetting
$\mathcal{R}=\{x\in\mathcal{X}\colon\text{there exists }(\underline{x},\underline{u})\in\mathcal{S}\text{ with }\underline{x}_{0}=x\}$
and point out that $\psi_{n}(\mathcal{R}\times\underline{\mathcal{U}})=\mathcal{R}$ for all $n\in\mathbb{N}$.
@@
The following implications and equivalences hold by definition.
From the given definition, it may not be immediately apparent why the IFP indeed conceptualizes ‘input forgetting’. To see how the IFP is motivated, consider $x^{1},x^{2}\in\mathcal{R}$. Take solutions $(\underline{x}^{1},\underline{u}^{1}),(\underline{x}^{2},\underline{u}^{2})\in\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}^{-}$ with $\underline{x}_{0}^{1}=x^{1}$ and $\underline{x}_{0}^{2}=x^{2}$. Then, $\psi_{n}(x^{i},\underline{u})$ is exactly the time zero component of a solution for the input $\underline{u}^{i,n}:=(\dots,\underline{u}_{-1}^{i},\underline{u}_{0}^{i},\underline{u}_{1},\dots,\underline{u}_{n})$. The $n$ most recent inputs in the sequences $\underline{u}^{1,n}$ and $\underline{u}^{2,n}$ are the same. Thus, although the IFP is defined through states $x^{1},x^{2}\in\mathcal{R}$, it really considers solutions of different input sequences whose difference lies further and further in the past as $n$ grows; see also Figure˜1.
Figure 1: Illustrations of the IFP (top), SFP (left), and s-SFP (right).
@@ 2.4  Attraction
We hinted in the introduction that one of the overarching principles is that the echo states form the dynamical attractor of the RNN to which the states converge. By the invariance $\psi_{n}(\mathcal{R}\times\underline{\mathcal{U}})=\mathcal{R}$, the SFP yields for all $\underline{u}\in\underline{\mathcal{U}}$ and $x\in\mathcal{X}$ [^2] Here, $\mathrm{dist}_{\mathcal{X}}(x,A)=\inf_{y\in A}d_{\mathcal{X}}(x,y)$ for a point $x\in\mathcal{X}$ and a subset $A\subseteq\mathcal{X}$.
@@
$\limsup_{n\rightarrow\infty}\mathrm{dist}_{\mathcal{X}}(\psi_{n}(x,\underline{u}),\mathcal{R})\leq\limsup_{n\rightarrow\infty}d_{\mathcal{X}}(\psi_{n}(x,\underline{u}),\psi_{n}(x^{\prime},\underline{u}))=0,$
where $x^{\prime}$ is an arbitrary element in $\mathcal{R}$. Likewise, the s-SFP yields
$\limsup_{n\rightarrow\infty}\mathrm{dist}_{\mathcal{X}}(\psi_{n}(x,T^{n}(\underline{u})),\mathcal{R})\leq\limsup_{n\rightarrow\infty}d_{\mathcal{X}}(\psi_{n}(x,T^{n}(\underline{u})),\psi_{n}(x^{\prime},T^{n}(\underline{u})))=0.$
Although both the SFP and the s-SFP imply attraction of $\mathcal{R}$, we now argue that the SFP captures forward attraction and the s-SFP pullback attraction. To make this claim rigorous, let us specify the precise dynamical system. Despite being intrinsically non-autonomous, it was proposed in [47] to model the dynamics autonomously on an extended sequence space. To this end, consider the map $\varphi\colon\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}\rightarrow\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}$ given by $\varphi(\underline{x},\underline{u})=((\underline{x},f(\underline{x}_{0},\underline{u}_{1})),\sigma(\underline{u}))$, whose iterates can be seen as extensions of the maps $\psi_{n}$. Concretely, $\psi_{n}\circ(p_{0}\times\pi)=p_{0}\circ\varphi^{n}$, where $p_{0}\colon\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}\rightarrow\mathcal{X}$ and $\pi\colon\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}\rightarrow\underline{\mathcal{U}}$ are the projections $p_{0}(\underline{x},\underline{u})=\underline{x}_{0}$ and $\pi(\underline{x},\underline{u})=\underline{u}$. It has been shown in [47] that the global attractor $\mathcal{A}:=\bigcap_{n\in\mathbb{N}}\varphi^{n}(\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}})$ of $\varphi$ is exactly equal to the inverse image $\tau^{-1}(\mathcal{S})$ of the set of solutions under the truncation $\tau\colon\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}\rightarrow\underline{\mathcal{X}}^{-}\times\underline{\mathcal{U}}^{-}$ extended to the product space. In particular, the set of reachable states $\mathcal{R}=p_{0}(\mathcal{A})$ is the $p_{0}$-projection of the attractor $\mathcal{A}$. The fibers $\mathcal{A}\cap\pi^{-1}(\underline{u})$ of the attractor capture pullback attraction, and the fibers $\mathcal{A}\cap\pi^{-1}(\sigma^{n}(\underline{u}))$ capture forward attraction. The relation $\psi_{n}\circ(p_{0}\times\pi)=p_{0}\circ\varphi^{n}$ implies that $\psi_{n}(x,\underline{u})$ belongs to the $p_{0}$-projection of the fiber $\pi^{-1}(\sigma^{n}(\underline{u}))$, and $\psi_{n}(x,T^{n}(\underline{u}))$ belongs to the $p_{0}$-projection of the fiber $\pi^{-1}(\underline{u})$, which together with the equality $\mathcal{R}=p_{0}(\mathcal{A})$ finally explains the association of the SFP with forward and of the s-SFP with pullback attraction. [^3] One can also consider the attractor $\mathcal{A}_{0}:=\bigcap_{n\in\mathbb{N}}\psi_{n}(\mathcal{X}\times\underline{\mathcal{U}})$ of the non-autonomous dynamics of $\psi_{n}$. But since the inclusion $p_{0}(\mathcal{A})\subseteq\mathcal{A}_{0}$ is strict in general, the attractor $\mathcal{A}$ gives a more detailed picture.
@@
2.5  Main result
In our main result, stated below, we extend the diagram in Remark˜2.2 with additional implication arrows under compactness assumptions. It generalizes [23, Proposition 1], [37, Theorem 1], and [30, Remark 1], and provides alternative and shorter proofs.
Theorem 2.4.
Consider the diagram below. If $\underline{\mathcal{X}}^{-}$ is compact, then the implication arrows indicated in blue are valid; and if $\underline{\mathcal{X}}^{-}$ and $\underline{\mathcal{U}}$ are compact, then the implication arrows indicated in green are valid. Furthermore, we will see counterexamples to the implication arrows indicated in purple.
@@ 2.6  Technical details
2.7  The significance of forgetting states
The SFP bears particular significance for applications. Suppose we are trying to learn an unknown invertible dynamical system as in Remark˜2.3, that is, the inputs are of the form $\underline{u}_{t}=\omega(\phi^{t}(p))$, and at the core of our learning method is the state-space system $f$. For example, $f$ could be an echo state network or the update map of a reservoir computer [23]. On an abstract level, successful learning manifests in the existence of a semi-conjugacy $\zeta\colon\mathcal{M}\rightarrow\mathcal{X}$, (meaning $f(\zeta(p),\omega(\phi(p)))=\zeta(\phi(p))$ for all $p\in\mathcal{M}$) that, moreover, is an embedding. Such a semi-conjugacy is called a generalized synchronization (GS) [31, 34]. A prerequisite for the existence of a GS is the ESP of the system. If the GS is an embedding, then it is, in fact, a homeomorphism from $\mathcal{M}$ to $\mathcal{R}$ [47] that, under certain hypotheses, becomes differentiable [17]. A trajectory $(\phi^{t}(p))_{t\in\mathbb{Z}_{-}}$ is synchronized through $\zeta$ with the unique solution $\underline{x}$ of the state-space system for the input $\underline{u}=(\omega(\phi^{t}(p)))_{t\in\mathbb{Z}_{-}}$. If we have access to the state $\underline{x}_{t}$ at time $t$, then applying the transformation $h:=\omega\circ\phi\circ\zeta^{-1}$ to $\underline{x}_{t}$ yields the next observation $\underline{u}_{t+1}$, that is, a perfect prediction. In practice, the map $h$ will be learned from data, giving rise to an approximate prediction $\hat{h}(\underline{x}_{t})\approx h(\underline{x}_{t})=\underline{u}_{t+1}$ [16].
This approximate prediction requires access to the state $\underline{x}_{t}$. The solution $\underline{x}$ is an object indexed in the infinite past. However, the actual implementation of $f$ begins at some point in time, say 0. Since we do not know the solution $\underline{x}$, we cannot initialize the system at $\underline{x}_{0}$, but instead rely on a random initialization or an educated guess $\underline{x}_{0}^{\prime}$ [16]. Since the initial state $\underline{x}_{0}^{\prime}$ contains imperfect information about the dynamics we are trying to learn, it is the SFP that guarantees that $\underline{x}_{t}^{\prime}$, the actual state of the implemented system at time $t$, will be close to $\underline{x}_{t}$ after a sufficient burn-in time after which $\hat{h}(\underline{x}_{t}^{\prime})\approx\hat{h}(\underline{x}_{t})\approx h(\underline{x}_{t})=\underline{u}_{t+1}$. In the language of [1], the SFP ensures that the implemented system is asymptotically synchronized.